Calculate SS variance and standard deviation for the following sample of n = 5 scores&
10, 3, 8, 0, 4 &
1
Expert's answer
2010-12-05T19:14:40-0500
The mean is
X = (10+3+8+0+4)/5 = 5
The variance is the average of the squared difference s from the mean:
σ 2 = (52+(-2)2+32+(-5)2+(-1)2)/5 = 12.8
The standard deviation equals to square root of the variance:
σ = √ (12.8) = 3.58